If the terminal side of angle ∅ is in first quadrant and tan ∅ is 1/√3. Find the values of the other trigonometric functions of ∅.

1
Expert's answer
2022-02-16T08:03:17-0500

Hypotenuse"=\sqrt{1+(√3)^2}=2"


"cot\>\theta=\sqrt{3}\>\>\>\>\>sin\>\theta=\frac{1}{2}"

"csc\>\theta=2\>\>\>\>\>cos\>\>\theta=\frac{√3}{2}"


"sec\>\theta=\frac{2}{√3}"